Lead thiocyanate, Pb(SCN)2, has a Ksp value of 2.00×10−5.

a) Calculate the molar solubility of lead thiocyanate in pure water. The molar solubility is the maximum amount of lead thiocyanate the solution can hold.

b) Calculate the molar solubility of lead thiocyanate in 0.500 M KSCN.
Answer #1

a)

Ksp = [Pb+2][SCN-]^2

Assume S = Pb+2, then SCN- = 2S

Ksp = (S)(2S)^2

2*10^-5 = (4s^3)

S = 0.0171 mol per liter (of Pb(SCN)2)

b)

if KSCN --> 0.5 M

then

Ksp = [Pb+2][SCN-]^2

2*10^-5 = (S)(0.50)^2

S = 8*10^-5 mol per liter (of Pb(SCN)2)